OpenStudy (anonymous):

What is the slope of a line with equation y=-3

OpenStudy (gorv):

equation of a line y=mx+c here m=slope

OpenStudy (gorv):

so compare with this your equation

Nnesha (nnesha):

slope equation is y = mx +b so you don't have any x so there is not any slope

OpenStudy (gorv):

term with x is not there means m=0 that is slope =0

Nnesha (nnesha):

slope is always with x on this equation y =mx + b